# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

### v1

@@ 6.1.2 Experiments and Results
$M_{\mathrm{agg}}$ achieves an F1 score of 30.3% on $D_{\mathrm{mc}}$, up from the baseline model $M_{\mathrm{mc}}$ ’s 27.3%, a notable 3-percentage point improvement. Additionally, $M_{\mathrm{agg}}$ is highly performant across the synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), surpassing $M_{\mathrm{mc}}$ by large margins on these. For a sense of what high performance means on these datasets, we reference the diagonal entries in rows for models $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, representing performance of these models on the same test data distribution as train: 83.9% for $D_{\mathrm{bs}}$, 75.4% for $D_{\mathrm{ss}}$, and 64.4% for $D_{\mathrm{r}}$. In light of these numbers, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it even performs worse on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. In contrast, $M_{\mathrm{agg}}$ is more robust, and retains much of the diagonal-entry performance on the synthetic datasets while also significantly improving on $D_{\mathrm{mc}}$. It is also better on the synthetic datasets than on $D_{\mathrm{mc}}$ by large margins (as it should be), unlike $M_{\mathrm{mc}}$. Lastly, the models trained solely on synthetic datasets perform poorly on $D_{\mathrm{mc}}$, as does $M_{\mathrm{mc}}$ on synthetic data, indicating that joint training on the aggregated dataset is essential for robust performance across both real and synthetic data.
@@ 6.2 Curriculum Learning (CL)
Another key observation from Table 4 is that performance of the $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, and $M_{\mathrm{mc}}$ models decreases in the order $M_{\mathrm{bs}}$ > $M_{\mathrm{ss}}$ > $M_{\mathrm{r}}$ > $M_{\mathrm{mc}}$, suggesting that the complexity/difficulty of examples increases in the order $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$ < $D_{\mathrm{mc}}$, with the most complex $D_{\mathrm{mc}}$ examples naturally appearing at the end. This observation raises the possibility of using this ordering to improve over the simple joint training approach (where data is randomly ordered) when training on the aggregated dataset, so as to achieve even better performance on the complex $D_{\mathrm{mc}}$. Curriculum Learning (CL) is a natural next step to explore in this direction. CL is a training strategy that sequences training data from simpler to more complex examples, similar to human learning. Introduced by Bengio et al. (2009), CL begins with easier examples, gradually increasing in difficulty. This progression allows the model to first learn basic patterns and representations, building a foundation before tackling more challenging cases, which can enhance generalization and improve performance on complex tasks.
@@ 6.2.1 CL Mechanics
A key prerequisite in CL is establishing difficulty metric(s) that quantify how difficult a training example is for a model to learn, allowing the data to be sorted accordingly during training. In this work, we explore three simple, natural, and intuitive difficulty metrics, along with their combinations (e.g., sorting first by one metric and then sorting (breaking ties) by another). The first metric is the dataset $D$ an example belongs to ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$). As mentioned before, the datasets can be ordered as $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$ < $D_{\mathrm{mc}}$ in terms of difficulty. Additionally, we consider the lengths of the action (output) and input sequences: the action (output) sequence length $L_{A}$ reflects the number of actions in the gold action sequence, while the input sequence length $L_{I}$ measures the number of tokens in the game history fed to the model. These are also intuitive as longer sequences are harder to generate as output and process as input.
@@
We make use of a simple CL algorithm, specifically using the Curriculum Arrangement algorithm within the CL framework proposed by Xu et al. (2020). Details of the method are provided in the appendix.
@@ 6.2.2 Experiments and Results
Overall experimental setup remains similar to the batch training one (Section 6.1.1). We experiment with each of the three difficulty metrics individually and the four possible combinations of them, making a total of seven metrics – $D$, $L_{A}$, $L_{I}$, $(D,L_{A})$, $(D,L_{I})$, $(L_{I},L_{A})$, and $(D,L_{I},L_{A})$ (ordering within a tuple denotes the sorting order). We also experiment with the CL algorithm’s hyperparameter, $N$, trying six values – 5, 10, 15, 20, 25, and 30. (This controls the number of buckets/difficulty levels – see appendix.) This results in a grid of $7\times 6=42$ experiments.
Table 5 presents the best model, $M_{\mathrm{agg+cl}}$, selected based on $D_{\mathrm{mc}}$ performance. For comparison, we also include the baseline $M_{\mathrm{mc}}$, and the aforementioned $M_{\mathrm{agg}}$ model (Table 4). On $D_{\mathrm{mc}}$, $M_{\mathrm{agg+cl}}$ achieves an F1 of 33.0%, a notable 2.7-point improvement over $M_{\mathrm{agg}}$, and an even more significant 5.7-point improvement over the baseline $M_{\mathrm{mc}}$. It also continues to be highly performant on the synthetic data similar to $M_{\mathrm{agg}}$. Thus, Curriculum Learning on the aggregated data provides a meaningful performance boost on the BAP task over simple joint training. Hereafter, we refer to $M_{\mathrm{agg+cl}}$ as the best model.
@@ 6.3 Transformers and LLMs (BERT)
To control the representation dimensionality and complexity of the decoder transformer, we add a simple feedforward projection layer atop BERT, compressing its output (as otherwise the dimensionality of the decoder transformer would be forced to be the same as the BERT encoder and make it overly complex than need be). The BERT encoder and transformer decoder are connected in standard encoder-decoder fashion but via the projection layer. BERT’s token embeddings are compressed by the projection layer, and the transformer decoder then attends to these representations at each decoding time-step, and the representation dimensionality of the decoder matches the output size of the projection layer. This also matches the embedding dimensionality for action vectors fed into the decoder. Positional embeddings are also used for the decoder input. The output of the decoder’s last layer conditions the CNN-based action predictor (replacing the GRU hidden state from the baseline model).
@@ 6.3.1 Experiments and Results
We conduct similar experiments with this model like with the GRU-based model, i.e., training on $D_{\mathrm{mc}}$ only, joint training on the aggregated dataset, and CL. The experimental setup and the CL experiments are same as before (Sections 6.1.1 and 6.2.2). We tested various sizes of BERT models (Turc et al., 2019), keeping BERT frozen, and found that $\text{BERT}_{\text{BASE}}$ (110.1M parameters) performed best. In the transformer decoder, five layers yielded optimal results, and a projection layer output size of 256 proved most effective. For tractability of experimentation, we only experimented with the number of transformer layers and the projection layer’s output size, keeping other architectural hyperparameters identical to the GRU-based models. Details specific to training regime, corresponding hyperparameters, etc. can be found in the appendix. Results are shown in Table 6. (Note: We also experimented with finetuning BERT, but found the frozen scenario to perform relatively better. However, as with the frozen case below, synthetic data and CL improved performance in the finetuning scenario as well, demonstrating the robustness of our approach. Results are provided in the appendix.)
@@
  • For tractability of experimentation, we only experimented with the number of transformer layers and the projection layer’s output size, keeping other architectural hyperparameters identical to the GRU-based models.
@@ 8.2 Evaluation on synthetic data
Going forward, evaluation on the synthetic data will also be important for making more meaningful progress on the BAP task. First, the more structured and simpler nature of the synthetic data allows us to assess basic model competencies. Second, as noted in Section 6.1.2, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it even performs worse on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. Evaluating on the synthetic data will help build models that do not show such counterintuitive behavior (Section 8.3 provides further evidence supporting this argument), and complements evaluation on $D_{\mathrm{mc}}$. In fact, future models should aim to surpass previous performance on $D_{\mathrm{mc}}$ while also achieving strong results on synthetic data, ideally improving on both.
@@ 8.3 Concurrent work
As mentioned in Section 1, modeling is not the primary focus of our work. Instead, we focus on the BAP task as a whole, addressing key challenges in data and evaluation to propose BAP v2, an upgraded version of the task itself. We also illustrate the impact of this data on training a simple LLM and transformer-based model. This foundation enables future work to make more efficient and meaningful progress on the task, and sets the stage for more sophisticated models/LLMs. These works are therefore complementary to ours. A slightly more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided in the appendix, given some other relevant aspects of their work. We, still, however, evaluate Nebula using our new BAP v2 benchmark to further demonstrate the impact and robustness of our evaluation, and to provide a fairer assessment of its performance. We also evaluate it on our synthetic data to further strengthen our argument about the importance of such evaluation as discussed in Section 8.2. Future work can explore finetuning models such as Llama-3-8B (Nebula) and other SOTA LLMs using our synthetic data and training methodologies like CL.
@@
Similar to the 29.4% improvement we observed for our baseline model $M_{\mathrm{mc}}$ under the v2 evaluation on $D_{\mathrm{mc}}$, Nebula’s performance also increases from 39.2% to 50.1%, reflecting a 27.8% gain. This further underscores the impact and robustness of our v2 evaluation, offering a fairer reflection of Nebula’s performance. Detailed tables for the full v2 benchmark are provided in the appendix (Tables 12 and 13).
@@
Table 9 shows Nebula’s performance on both the synthetic data ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) and the original $D_{\mathrm{mc}}$ test sets. Evaluating Nebula on the synthetic data reveals behavior consistent with our baseline model $M_{\mathrm{mc}}$, as reported in Section 6.1.2 and further discussed in Section 8.2. It underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it even performs worse on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. This shows that models trained solely on $D_{\mathrm{mc}}$ may not be as robust, and reinforces our argument from Section 8.2 about the importance of synthetic data evaluation: it will help build models that don’t show such counter-intuitive behavior.
@@ 9.1 Task and Dataset
The MDC and BAP (Narayan-Chen, Jayannavar, and Hockenmaier, 2019; Jayannavar, Narayan-Chen, and Hockenmaier, 2020) continue to remain a challenging and relevant testbed for grounded language instruction following. We discuss a few more recent works to highlight this.
@@
This limitation is addressed by the more recent benchmark TEACh (Padmakumar et al., 2022), which includes free-form dialogue, similar to the MDC, and features over 3000 human-human interactive dialogues for completing household tasks in simulation and with object state changes. However, unlike the MDC, the instruction giver in TEACh is provided with a predefined plan of steps, while the Architect in the MDC independently devises a unique plan for constructing the target structure. This results in greater creativity and diversity in the MDC, with multiple possible plans for the same structure. Although the TEACh dataset is larger (#dialogs), its dialogues and utterances are shorter. Additionally, the MDC allows for greater abstraction in instructions, such as referring to substructures, individual blocks, or shapes, whereas TEACh focuses on well-defined objects in the environment. Similar to VLN and CVDN, referring expressions in TEACh refer to objects that exist in the world, but construction instructions like those in the MDC frequently refer to objects that need to be built.
Minecraft, in recent years, has garnered a lot of interest as an AI experimentation platform. Our previous work (Narayan-Chen, Jayannavar, and Hockenmaier, 2019) was among the first to use it to study grounded task-oriented dialogue, and introduced the MCBT and MDC. Facebook’s CraftAssist (Gray et al., 2019; Jernite et al., 2019; Szlam et al., 2019) is another such example, enabling two-way human-bot interactions where a human architect instructs an automated builder to build complex structures. Their data includes synthetically generated and crowdsourced instructions paired with logical tree structures consisting of action primitives, unlike the MDC, which features human-human dialogues with greater ambiguity, variety, and noisier Builder actions. MineDojo (Fan et al., 2022) focuses on creating versatile agents for diverse tasks via an internet-scale knowledge base, contrasting with tasks like IGLU (Mohanty et al., 2024; Kiseleva et al., 2022b, a) and MCBT, which prioritize grounded natural language dialogue and clarification for interactive agents.

### v2

@@ 6.1.2 Experiments and Results
$M_{\mathrm{agg}}$ achieves an F1 score of 30.3% on $D_{\mathrm{mc}}$, up from the baseline model $M_{\mathrm{mc}}$ ’s 27.3%, a notable 3-percentage point improvement. Additionally, $M_{\mathrm{agg}}$ is highly performant across the synthetic datasets ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$), surpassing $M_{\mathrm{mc}}$ by large margins on these. For a sense of what high performance means on these datasets, we reference the diagonal entries in rows for models $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, representing performance of these models on the same test data distribution as train: 83.9% for $D_{\mathrm{bs}}$, 75.4% for $D_{\mathrm{ss}}$, and 64.4% for $D_{\mathrm{r}}$. In light of these numbers, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. In contrast, $M_{\mathrm{agg}}$ is more robust, and retains much of the diagonal-entry performance on the synthetic datasets while also significantly improving on $D_{\mathrm{mc}}$. It is also better on the synthetic datasets than on $D_{\mathrm{mc}}$ by large margins (as it should be), unlike $M_{\mathrm{mc}}$. Lastly, the models trained solely on synthetic datasets perform poorly on $D_{\mathrm{mc}}$, as does $M_{\mathrm{mc}}$ on synthetic data, indicating that joint training on the aggregated dataset is essential for robust performance across both real and synthetic data.
@@ 6.2 Curriculum Learning (CL)
Another key observation from Table 4 is that performance of the $M_{\mathrm{bs}}$, $M_{\mathrm{ss}}$, $M_{\mathrm{r}}$, and $M_{\mathrm{mc}}$ models decreases in the order $M_{\mathrm{bs}}$ > $M_{\mathrm{ss}}$ > $M_{\mathrm{r}}$ > $M_{\mathrm{mc}}$, suggesting that the complexity/difficulty of examples increases in the order $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$ < $D_{\mathrm{mc}}$, with the most complex $D_{\mathrm{mc}}$ examples naturally appearing at the end. This observation raises the possibility of using this ordering to improve over the naive joint training approach (where data is randomly ordered) when training on the aggregated dataset, so as to achieve even better performance on the complex $D_{\mathrm{mc}}$. Curriculum Learning (CL) is a natural next step to explore in this direction. CL is a training strategy that sequences training data from simpler to more complex examples, similar to human learning. Introduced by Bengio et al. (2009), CL begins with easier examples, gradually increasing in difficulty. This progression allows the model to first learn basic patterns and representations, building a foundation before tackling more challenging cases, which can enhance generalization and improve performance on complex tasks.
@@ 6.2.1 CL Mechanics
A key prerequisite in CL is establishing difficulty metric(s) that quantify how difficult a training example is for a model to learn, allowing the data to be sorted accordingly during training. In this work, we explore three straightforward, natural, and intuitive difficulty metrics, along with their combinations (e.g., sorting first by one metric and then sorting (breaking ties) by another). The first metric is the dataset $D$ an example belongs to ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $D_{\mathrm{mc}}$). As mentioned before, the datasets can be ordered as $D_{\mathrm{bs}}$ < $D_{\mathrm{ss}}$ < $D_{\mathrm{r}}$ < $D_{\mathrm{mc}}$ in terms of difficulty. Additionally, we consider the lengths of the action (output) and input sequences: the action (output) sequence length $L_{A}$ reflects the number of actions in the gold action sequence, while the input sequence length $L_{I}$ measures the number of tokens in the game history fed to the model. These are also intuitive as longer sequences are harder to generate as output and process as input.
@@
We make use of a simple CL algorithm, specifically using the Curriculum Arrangement algorithm within the CL framework proposed by Xu et al. (2020). Details of the method are provided in the appendix (Section 12.1).
@@ 6.2.2 Experiments and Results
Overall experimental setup remains similar to the one used for joint training (Section 6.1.1). We experiment with each of the three difficulty metrics individually and the four possible combinations of them, making a total of seven metrics – $D$, $L_{A}$, $L_{I}$, $(D,L_{A})$, $(D,L_{I})$, $(L_{I},L_{A})$, and $(D,L_{I},L_{A})$ (ordering within a tuple denotes the sorting order). We also experiment with the CL algorithm’s parameter, $N$, trying six values – 5, 10, 15, 20, 25, and 30. (This controls the number of buckets/difficulty levels – see appendix (Section 12.1).) This results in a grid of $7\times 6=42$ experiments.
Table 5 presents the best model, $M_{\mathrm{agg+cl}}$, selected based on $D_{\mathrm{mc}}$ performance. For comparison, we also include the baseline $M_{\mathrm{mc}}$, and the aforementioned $M_{\mathrm{agg}}$ model (Table 4). On $D_{\mathrm{mc}}$, $M_{\mathrm{agg+cl}}$ achieves an F1 of 33.0%, a notable 2.7-point improvement over $M_{\mathrm{agg}}$, and an even more significant 5.7-point improvement over the baseline $M_{\mathrm{mc}}$. It also continues to be highly performant on the synthetic data similar to $M_{\mathrm{agg}}$. Thus, Curriculum Learning on the aggregated data provides a meaningful performance boost on the BAP task over naive joint training. Hereafter, we refer to $M_{\mathrm{agg+cl}}$ as the best model.
@@ 6.3 Transformers and LLMs (BERT)
To control the representation dimensionality and complexity of the decoder transformer, we add a feedforward projection layer atop BERT, compressing its output (as otherwise the dimensionality of the decoder transformer would be forced to be the same as the BERT encoder and make it overly complex than need be). The BERT encoder and transformer decoder are connected in standard encoder-decoder fashion but via the projection layer. BERT’s token embeddings are compressed by the projection layer, and the transformer decoder then attends to these representations at each decoding time-step, and the representation dimensionality of the decoder matches the output size of the projection layer. This also matches the embedding dimensionality for action vectors fed into the decoder. Positional embeddings are also used for the decoder input. The output of the decoder’s last layer conditions the CNN-based action predictor (replacing the GRU hidden state from the baseline model).
@@ 6.3.1 Experiments and Results
We conduct similar experiments with this model like with the GRU-based model, i.e., training on $D_{\mathrm{mc}}$ only, joint training on the aggregated dataset, and CL. The experimental setup and the CL experiments are the same as for the GRU scenario (Sections 6.1.1 and 6.2.2). Further details specific to the BERT scenario are provided in the appendix (Section 12.4). We tested various sizes of BERT models (Turc et al., 2019), and keeping BERT frozen versus finetuning it. We found that frozen $\text{BERT}_{\text{BASE}}$ (110.1M parameters) performed the best. In the transformer decoder, five layers yielded optimal results, and a projection layer output size of 256 proved most effective. Other hyperparameters are provided in the appendix (Section 12.4). For tractability of experimentation, we only experimented with the transformer’s and projection layer’s hyperparameters, keeping other architectural hyperparameters identical to the GRU-based models (e.g. those of the CNNs). Results are shown in Table 6. (Note: We report only the best-performing BERT-based model, which was from the frozen scenario as mentioned earlier. However, similar to the results for the frozen case discussed below, synthetic data, joint training, and CL also improved model performance in the finetuning scenario, validating our approach.)
@@
  • For tractability of experimentation, we only experimented with the transformer’s and projection layer’s hyperparameters, keeping other architectural hyperparameters identical to the GRU-based models (e.g. those of the CNNs).
@@ 8.2 Evaluation on synthetic data
Going forward, evaluation on the synthetic data will also be important for making more meaningful progress on the BAP task. First, the more structured and simpler nature of the synthetic data allows us to assess basic model competencies. Second, as noted in Section 6.1.2, the baseline $M_{\mathrm{mc}}$ underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. Evaluating on the synthetic data will help build models that do not show such counterintuitive behavior (Section 8.3 provides further evidence supporting this argument), and complements evaluation on $D_{\mathrm{mc}}$. In fact, future models should aim to surpass previous performance on $D_{\mathrm{mc}}$ while also achieving strong results on synthetic data, ideally improving on both.
@@ 8.3 Concurrent work
As mentioned in Section 1, modeling is not the primary focus of our work. Instead, we focus on the BAP task as a whole, addressing key challenges in data and evaluation to propose BAP v2, an upgraded version of the task itself. We also illustrate the impact of this data on training a simple LLM and transformer-based model. This foundation enables future work to make more efficient and meaningful progress on the task, and sets the stage for more sophisticated models/LLMs. These works are therefore complementary to ours. A more detailed comparison to Chaturvedi, Thompson, and Asher (2024) is provided in the appendix (Section 15), given some other relevant aspects of their work.
Here, we evaluate Nebula using the v2 benchmark to further demonstrate the impact and robustness of our evaluation while providing a fairer assessment of its performance. We observe a 27.8% gain in performance (compared to the legacy evaluation), similar to the gain for our baseline $M_{\mathrm{mc}}$. We also discuss how the v2 evaluation now enables more accurate benchmarking of SOTA LLMs on the BAP task and highlights the significant room for improvement, further motivating the task itself. Additionally, it helps us confirm that spatial reasoning remains the key bottleneck for high performance, motivating future efforts leveraging our data and methods, which have shown effectiveness in addressing this challenge. We also evaluate Nebula on our synthetic data and find that, like the baseline $M_{\mathrm{mc}}$ (and the BERT-based $M_{\mathrm{mc}}^{\prime}$), it underperforms significantly despite being trained on the far more complex $D_{\mathrm{mc}}$. This suggests that models (including SOTA LLMs) trained solely on $D_{\mathrm{mc}}$ may not be as robust, reinforcing our argument of Section 8.2 about the importance of the synthetic data evaluation. Future work should explore finetuning models such as Llama-3-8B (Nebula) and newer LLMs using our synthetic data and training methodologies like CL.
@@
Similar to the 29.4% improvement we observed for our baseline model $M_{\mathrm{mc}}$ under the v2 evaluation on $D_{\mathrm{mc}}$, Nebula’s performance also increases from 39.2% to 50.1%, reflecting a 27.8% gain. This further underscores the impact and robustness of our v2 evaluation, offering a fairer reflection of Nebula’s performance. With this fairer evaluation now possible, we can more accurately gauge current performance on the BAP task and the remaining gap for improvement. While the 50.1% score marks improved performance, it also underscores that even with SOTA LLMs, significant room for improvement remains, thus reinforcing the task’s challenging nature and its importance for future research directions. Detailed tables for the full v2 benchmark are provided in the appendix (Tables 13 and 14). The observed trends generally align with those seen in our models such as the model’s relative strengths and weaknesses. They also shed some light on key areas for future improvement. For example, spatial reasoning (Location F1) is the key bottleneck for high performance on this task as noted in Section 4.4, and this continues to remain the case. Therefore, as demonstrated with $M_{\mathrm{mc}}$, our synthetic data and training methodologies can potentially help address this challenge, motivating future work.
@@
Table 9 shows Nebula’s performance (micro F1) on both the synthetic data ($D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$) and the original $D_{\mathrm{mc}}$ test sets. Evaluating Nebula on the synthetic data reveals behavior consistent with our baseline model $M_{\mathrm{mc}}$ (and the BERT-based $M_{\mathrm{mc}}^{\prime}$), as reported in Section 6.1.2 and further discussed in Section 8.2 – it underperforms significantly on the synthetic datasets, despite being trained on the far more complex $D_{\mathrm{mc}}$ data; it also performs lower on datasets $D_{\mathrm{ss}}$ and $D_{\mathrm{r}}$ than on $D_{\mathrm{mc}}$. This shows that models (including SOTA LLMs) trained solely on $D_{\mathrm{mc}}$ may not be as robust, and reinforces our argument from Section 8.2 about the importance of the synthetic data evaluation: it will help build models that don’t show such counter-intuitive behavior.
@@ 9.1 Task and Dataset
We discuss related work to illustrate that the MDC and BAP remain a challenging and relevant testbed for grounded language instruction following.
@@
This limitation is addressed by the more recent benchmark TEACh (Padmakumar et al., 2022), which includes free-form dialogue, similar to the MDC, and features over 3000 human-human interactive dialogues for completing household tasks in simulation and with object state changes. However, unlike the MDC, the instruction giver in TEACh is provided with a predefined plan of steps, while the Architect in the MDC independently devises a unique plan for constructing the target structure. This results in greater creativity and diversity in the MDC, with multiple possible plans for the same structure. Although the TEACh dataset is larger (by number of #dialogs), its dialogues and utterances are shorter. Additionally, the MDC allows for greater abstraction in instructions, such as referring to substructures, individual blocks, or shapes, whereas TEACh focuses on well-defined objects in the environment. Similar to VLN and CVDN, referring expressions in TEACh refer to objects that exist in the world, but construction instructions like those in the MDC frequently refer to objects that need to be built.
Minecraft has garnered a lot of interest as an AI experimentation platform in recent years. Our previous work (Narayan-Chen, Jayannavar, and Hockenmaier, 2019) was among the first to use it to study grounded task-oriented dialogue, and introduced the MCBT and MDC. Facebook’s CraftAssist (Gray et al., 2019; Jernite et al., 2019; Szlam et al., 2019) is another such example, enabling two-way human-bot interactions where a human architect instructs an automated builder to build complex structures. Their data includes synthetically generated and crowdsourced instructions paired with logical tree structures consisting of action primitives, unlike the MDC, which features human-human dialogues with greater ambiguity, variety, and noisier Builder actions. MineDojo (Fan et al., 2022) focuses on creating versatile agents for diverse tasks via an internet-scale knowledge base, contrasting with tasks like IGLU (Mohanty et al., 2024; Kiseleva et al., 2022b, a) and MCBT, which prioritize grounded natural language dialogue and clarification for interactive agents.